The Extraordinary Manufacturing Company has established that the relationship between sales price for one of its products and the quantity sold per month is approximately 3D=1680-39p units. The fixed cost is P1500 per month, the variable cost is P45 per unit produced.


  • What number of units should be produced and sold per month to maximize net profit? What is the maximum profit?
  • What is the production range of profitable business for the company?
1
Expert's answer
2021-08-12T05:09:37-0400

Given

"3D=1680-3P"

Fixed cost=1500

Variable cost=45

Rearrange

"3D=1680-3D\\3P=1680-3D\\P=560-D"

a)

Profit maximizing level

"MR=MC\\TR=P×D\\TR=(560-D)D^2\\MR=\frac{dTR}{dD}=560-2D\\TC=1500+45D\\MC=\frac{dTC}{dD}=45"

Now MR=MC

"560-2D=45\\560-45=2D\\D=257.5\\P=560-D\\P=560-257.5=302.5"


Profit=TR-TC

"TR=560D-D^2\\=560(257.5)-(257.5×257.5)\\=144200-66306.25\\=77893.75"

"TC=1500+45D\\=1500+45(257.5)\\=1500+11.587.5\\=13087.5"

Profit"=77893.75-13087.5=64806.25"


b)

"Range=257.5-0.5=257\\257+0.5=258\\257-258"